On February 12 (day 43), the NOAA-15 HIRS instrument began showing
affects from increased Filter Motor Current. The Filter Motor Current
began to rise on orbit 19498 but didn't seriously affect Sync until
orbit 19503 when the sync went below 0.999. Sync is currently 70% (0.7)
throughout orbits. Channel 7 space view went into saturation at the
time of sync loss causing complete data loss for ch7 processing.